Working together, two pumps can drain a certain pool in 3 hours. If it takes the older pump 9 hours to drain the pool by itself, how long will it take the newer pump to drain the pool on its own?
Answer by TimothyLamb(4379)

flow rate = r = v/t
t = v/r
---
v = volume of pool
x = rate of old pump
y = rate of new pump
---
x + y = v/t
x + y = v/3
x = v/9
---
x + y = v/3
v/9 + y = v/3
y = v/3 - v/9
y = 3v/9 - v/9
y = 2v/9
y = (2/2)v/(9/2)
y = v/(9/2)
answer:
(9/2) = 4.5 hours
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php